Find the critical values, z-scores, for the following confidence intervals by showing a graph and respective areas:

a) 80% z-scores b) 95% t-scores. n = 81
Answer #1

a)
Given CI level is 80%, hence α = 1 - 0.8 = 0.2
α/2 = 0.2/2 = 0.1, Zc = Z(α/2) = 1.28

b)
Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 1.99